Find a quadratic model for the set of values (-2, -20)(0,-4) (4, -20) if u can, can u show your work please thnx so much, trying to see if i did it wright.
Answer by ankor@dixie-net.com(22740) 
 Find a quadratic model for the set of values (-2, -20)(0,-4) (4, -20)
:
Develop an equation from each set of points using the form: ax^2 + bx + c = y
Find a, b, c. The 2nd set of values, 0.-4, tells us c=-4 (x=0)
:
x=-2; y=-20
4a - 2b - 4 = -20
4a - 2b = -20 + 4
4a - 2b = -16
:
x=4; y=-20
16a + 4b - 4 = -20
16a + 4b = -20 + 4
16a + 4b = -16
simplify, divide by 2
8a + 2b = -8
:
Use elimination, add the two equations
4a - 2b = -16
8a + 2b = -8
--------------adding eliminates b, find a
12a = -24
a = -24/12
a = -2
:
Find b using the simplified equatin 8a + 2b = -8
8(-2) + 2b = -6
-16 + 2b = -8
2b = -8 + 16
2b = 8
b = 4
:
The equation y = -2x^2 + 4x - 4
:
:
You can confirm this for yourself, replace the x with given values and see if they equal the given y
| |
|